# Communication-Efficient Distributed Deep Learning via Federated Dynamic Averaging

### v1

@@ Abstract.
Driven by the ever-growing volume and decentralized nature of data, coupled with the escalating size of modern models, distributed deep learning (DDL) has been entrenched as the preferred paradigm for training. However, frequent synchronization of DL models, encompassing millions to many billions of parameters, creates a communication bottleneck, severely hindering scalability. Worse yet, DDL algorithms typically waste valuable bandwidth, and make themselves less practical in bandwidth-constrained federated settings, by relying on overly simplistic, periodic, and rigid synchronization schedules. To address these shortcomings, we propose Federated Dynamic Averaging (FDA), a communication-efficient DDL strategy that dynamically triggers synchronization based on the value of the model variance. Through extensive experiments across a wide range of learning tasks we demonstrate that FDA reduces communication cost by orders of magnitude, compared to both traditional and cutting-edge communication-efficient algorithms. Remarkably, FDA achieves this without sacrificing convergence speed — in stark contrast to the trade-offs encountered in the field. Additionally, we show that FDA maintains robust performance across diverse data heterogeneity settings.
@@ 1. Introduction
Deep Learning (DL) stands at the forefront of the most significant technological advancements in recent times, with Deep Neural Networks (DNNs) achieving remarkable success in fields such as computer vision (Kolesnikov et al., 2020), natural language processing (Devlin et al., 2019; Brown et al., 2020) and speech recognition (Amodei et al., 2015). These rapid advancements coincide with the big data era, marked by an unprecedented scale of training datasets (Sun et al., 2017; Russakovsky et al., 2015). These datasets are not only growing in size but are often physically distributed and cannot be easily centralized due to business considerations, privacy concerns, bandwidth limitations, and data sovereignty laws (Kairouz et al., 2019; McMahan et al., 2016; Konečný et al., 2016). In response, Distributed Deep Learning (DDL) has emerged as an alternative paradigm to the traditional centralized approach (Zinkevich et al., 2010; Peteiro-Barral and Guijarro-Berdiñas, 2012; Chilimbi et al., 2014), offering efficient handling of large-scale data across multiple worker-nodes, enhancing the speed of training DL models and paving the way for more scalable and resilient DL applications (Moritz et al., 2016; Abadi et al., 2015; Li et al., 2020).
@@
Most DDL methods are iterative, where, in each iteration, some amount of local training is followed by synchronization of the local model with the global one. The predominant method, based on the bulk synchronous parallel (BSP) approach (Valiant, 1990), is to average the local model updates and then apply the average update to each local model (Zinkevich et al., 2010; Dekel et al., 2010). Less synchronized variants have also been proposed, to ameliorate the effect of straggler workers (Niu et al., 2011; Cui et al., 2014; Gupta et al., 2016; Mitliagkas et al., 2016; Dutta et al., 2018; Ho et al., 2013) but sometimes compromise convergence speed and model quality.
@@
A critical challenge inherent in the traditional techniques, especially in federated DL settings, where models are huge and worker interconnections are slow, is the communication bottleneck, restricting system scalability (Shi and Chu, 2017; Tang et al., 2023). This leads to a low computation-to-communication ratio (Shi et al., 2021; Lin et al., 2017). Addressing the communication overhead to expedite DDL algorithms has been a focal point of research for many years; speeding-up SGD is arguably among the most impactful and transformative problems in machine learning (Wang and Joshi, 2021).
@@
The most direct method to alleviate the communication burden is to reduce the frequency of communication rounds. Local-SGD is the prime example of this approach. It allows workers to perform $\tau$ local update steps on their models before aggregating them, as opposed to averaging the updates in every step (Zhang et al., 2016; Yu et al., 2018; Haddadpour et al., 2019). Although Local-SGD is effective in reducing communication while maintaining similar accuracy levels (Smith et al., 2016; Wang and Joshi, 2021), determining the optimal value of $\tau$ presents a significant challenge, with only a handful of studies offering theoretical insights into its influence on convergence (Wang and Joshi, 2021; Stich, 2019; Zhou and Cong, 2017; Yu et al., 2018).
@@
To further reduce communication costs of Local-SGD, more sophisticated communication strategies introduce varying sequences of local update steps $\{\tau_{0},...,\tau_{R}\}$, instead of a fixed $\tau$. In (Wang and Joshi, 2018), in order to minimize convergence error with respect to wall-time, the authors proposed a decreasing sequence of local update steps. Conversely, the focus in (Haddadpour et al., 2019) was on reducing the number of communication rounds for a fixed number of model updates and prescribed an increasing sequence. These contrasting approaches underscore the multifaceted nature of communication strategies in distributed deep learning, highlighting not only the absence of a one-size-fits-all solution but also the growing need for dynamic, context-aware strategies that can continuously adapt to the specific intricacies of the learning task.
@@
Main Contributions. Our work addresses critical efficiency challenges in DDL, particularly in communication-constraint environments, such as the ones encountered in Federated Learning (FL) (Kairouz et al., 2019) applications. We introduce Federated Dynamic Averaging (FDA), a novel, adaptive distributed deep learning strategy, massively improving communication efficiency over previous work. Our contributions are as follows:
@@
  • FDA dynamically decides to synchronize local nodes when model variance across nodes exceeds a threshold. This strategy drastically reduces communication requirements, while preserving cohesive progress towards the shared training objective.
  • We evaluate and compare FDA with other DDL algorithms through a comprehensive suite of experiments with diverse datasets, models, and tasks. This extensive approach—requiring over $200$K GPU hours—compensates for the stochasticity associated with training DNNs, thereby enabling us to draw more definitive conclusions.
@@
  • We demonstrate that FDA outperforms traditional and contemporary FL algorithms by orders of magnitude in communication, while maintaining equivalent model performance. Furthermore, it effectively balances the competing demands of reducing communication and computation, providing greatly improved trade-offs.
  • We show that FDA remains robust and efficient under various data distributions (Non-IID settings), maintaining comparable performance to the IID case. Furthermore, we demonstrate that FDA leads to models with excellent generalization, inherently mitigating overfitting exhibited by previous techniques.
@@ 2. Related Work
where $K$ is the number of workers and $F_{k}(\mathbf{w})\triangleq\mathbb{E}_{\zeta_{k}\sim\mathcal{D}_{k}}\left[\ell(\mathbf{w};\zeta_{k})\right]$ is the local objective function for worker $k$. Function $\ell(\mathbf{w};\zeta_{k})$ represents the loss for data sample $\zeta_{k}$ given model $w$.
@@
Communication efficient Local-SGD. The work in (Li et al., 2021) decomposes each round into two phases. In the first phase, each worker runs Local-SGD with $\tau=I_{1}$, while the second phase run $I_{2}$ steps with $\tau=1$(Li et al., 2021) proposes to exponentially decay $I_{1}$ every $M$ rounds. In the heterogeneous setting, the work in (Qin et al., 2023), by analysing the convergence rate, proposes an increasing sequence of local update steps for strongly-convex local objectives, and suggests fixed local steps for convex, non-convex local objectives. The study in (Yu and Jin, 2019) dynamically increases batch sizes to reduce communication rounds, maintaining the same convergence rate as SSP-SGD. However, the large-batch approach leads to poor generalization (Lin et al., 2018; Hoffer et al., 2018; Shallue et al., 2018), a challenge addressed by the post-local SGD method (Lin et al., 2018), which divides training into two phases: BSP-SGD followed by Local-SGD with a fixed number of steps. In the Lazily Aggregated Algorithm (LAG) (Chen et al., 2018), a different approach was taken, using only new gradients from some selected workers and reusing the outdated gradients from the rest, which essentially skips communication rounds.
@@
Federated Averaging (FedAvg) (McMahan et al., 2016) is another representative communication efficient Local-SGD algorithm, which is a pivotal method in Federated Learning (FL) (Kairouz et al., 2019). In the FL setting with edge computing systems, the work in (Wang et al., 2019) tries to find the optimal synchronization period $\tau$ subject to local computation and aggregation constraints. Recently (Mills et al., 2023), in the FL setting with the assumption of strongly-convex objectives, by analysing the balance between fast convergence and higher-round completion rate, a decaying local update step scheme emerged.
@@
Accelerating convergence. An indirect, yet highly effective, way to mitigate the communication burden in DDL, is to speed up convergence. Consequently, recent works have deployed accelerated versions of SGD to the distributed setting. Specifically, FedAdam (Reddi et al., 2021) extends Adam (Kingma and Ba, 2017) and FedAvgM (Hsu et al., 2019) extends SGD with momentum (SGD-M) (Sutskever et al., 2013a). Recently, Mime (Karimireddy et al., 2021) provides a framework to adapt arbitrary centralized optimization algorithms to the FL setting. However, these methods still suffer from the model divergence problem, particularly in heterogeneous settings. When solving (1), the disparity between each worker’s optimal solution $\mathbf{w}_{k}^{*}$ for their objective $F_{k}$, and the global optimum $\mathbf{w}^{*}$ for $F$, can potentially cause worker models to diverge (drift) towards their disparate minima (Karimireddy et al., 2019; Reddi et al., 2021). The result is slow and unstable convergence with significant communication overhead. To address this problem, the SCAFFOLD algorithm (Karimireddy et al., 2019) used control-variates (in the same spirit to SVRG (Johnson and Zhang, 2013)), with significant speed-up. FedProx (Sahu et al., 2018) re-parametarized FedAvg (McMahan et al., 2016) by adding $L^{2}$ regularization in the workers’ objectives to be near the global model. Lastly, FedDyn (Acar et al., 2021) improved upon these ideas with a dynamic regularizer making sure that if local models converge to a consensus, this consensus point aligns with the stationary point of the global objective function.
@@
Compression. To reduce communication overhead in DDL, significant efforts have been directed towards minimizing message sizes. Key strategies include sparsification, where only crucial components of information are transmitted, as explored in the works (Ström, 2015; Aji and Heafield, 2017; Alistarh et al., 2018), and quantization techniques, which involve transmitting only quantized gradients, as detailed in studies like (Rabbat and Nowak, 2005; Shlezinger et al., 2020). Notably, these compression techniques can be effectively combined with Local-SGD methods to enhance communication-efficiency further. A prime example is Qsparse-local-SGD (Basu et al., 2019), which integrates aggressive sparsification and quantization with Local-SGD, achieving substantial communication savings to reach specific target accuracies. A comprehensive survey of compression techniques in DDL can be found in (Wang et al., 2023).
@@ 3. Federated Dynamic Averaging
Notation. At each time instance $t$, each worker $k$ independently maintains its own vector of model parameters, [^1] The terms “model” and ”model parameters” are used interchangeably, as is common in the literature. denoted as $\mathbf{w}_{t}^{(k)}\in\mathbb{R}^{d}$. Let $\mathbf{\Delta w}^{(k)}_{t}$ be the update vector (Goodfellow et al., 2016) computed by some optimization algorithm (e.g., SGD, Adam) encompassing the learning rate and relevant gradients. In a distributed learning step $t$, each worker $k$ applies the following update:
(2) $\mathbf{w}^{(k)}_{t+1}=\mathbf{w}^{(k)}_{t}+\mathbf{\Delta w}^{(k)}_{t}$
Let $\overline{\mathbf{w}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{t}^{(k)}$ represent the average model at time $t$ and $\overline{\mathbf{w}}_{t_{0}}$ the initial model shared among the workers at the start of the current communication round. Moreover, we introduce the drift vector $\mathbf{u}_{t}^{(k)}=\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t_{0}}$, that is, the change to the local model of the $k$-th worker at time $t$ since the beginning of the current round at time $t_{0}$, and $\overline{\mathbf{u}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{u}_{t}^{(k)}=\overline{\mathbf{w}}_{t}-\overline{\mathbf{w}}_{t_{0}}$ as the average drift.
Model Variance. The model variance quantifies the dispersion or spread of the worker models around the average model. At time $t$, it is defined as:
@@
$\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t}\right\|_{2}^{2}=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{w}_{t}^{(k)}-\mathbf{w}_{t_{0}}+\mathbf{w}_{t_{0}}-\overline{\mathbf{w}}_{t}\right\|_{2}^{2}$
@@
(3) $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}-\overline{\mathbf{u}}_{t}\right\|_{2}^{2}=...=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
@@
This measure provides insight into how closely aligned the workers’ models are at any given time. High variance indicates that the models are widely spread out, essentially drifting apart, leading to a lack of cohesion in the aggregated model. Conversely, a moderate-low variance suggests that the workers’ models are closely aligned, working collectively towards the shared objective. The variance plays a critical role in our approach, as it helps us gauge the state of the training process making informed decisions about round termination.
@@
Round Invariant. We introduce the Round Invariant (RI), a condition that bounds model variance below a threshold $\Theta$:
@@
Our algorithm continuously monitors the model variance to maintain the RI. If the variance estimate exceeds the threshold $\Theta$ at any point $t$, the training round immediately terminates. This triggers synchronization, where all local model weights $\mathbf{w}^{(k)}_{t}$ are averaged and consolidated into the new global model $\overline{\mathbf{w}}_{t}$.
Monitoring the RI. At this point, our focus shifts to devising a method to monitor the RI, as defined in (4). We begin by defining the local state for each worker $k$, denoted by $S_{k}(t)\in\mathbb{R}^{p}$, which encapsulates information from worker $k$ at time $t$; it is updated arbitrarily. Next, we introduce the global state, $S(t)\in\mathbb{R}^{p}$, which represents the collective state of the distributed system at time $t$.
(5) $S(t)=\frac{1}{K}\sum_{k=1}^{K}S_{k}(t)$
Lastly, we define a special non-linear function $H:\mathbb{R}^{p}\to\mathbb{R}$, with the following property:

### v2

@@ Abstract.
Driven by the ever-growing volume and decentralized nature of data, coupled with the need to harness this data and generate knowledge from it, has led to the extensive use of distributed deep learning (DDL) techniques for training. These techniques rely on local training that is performed at the distributed nodes based on locally collected data, followed by a periodic synchronization process that combines these models to create a global model. However, frequent synchronization of DL models, encompassing millions to many billions of parameters, creates a communication bottleneck, severely hindering scalability. Worse yet, DDL algorithms typically waste valuable bandwidth, and make themselves less practical in bandwidth-constrained federated settings, by relying on overly simplistic, periodic, and rigid synchronization schedules. These drawbacks also have a direct impact on the time required for the training process, necessitating excessive time for data communication. To address these shortcomings, we propose Federated Dynamic Averaging (FDA), a communication-efficient DDL strategy that dynamically triggers synchronization based on the value of the model variance. In essence, the costly synchronization step is triggered only if the local models, which are initialized from a common global model after each synchronization, have significantly diverged. This decision is facilitated by the communication of a small local state from each distributed node/worker. Through extensive experiments across a wide range of learning tasks we demonstrate that FDA reduces communication cost by orders of magnitude, compared to both traditional and cutting-edge communication-efficient algorithms. Additionally, we show that FDA maintains robust performance across diverse data heterogeneity settings.
Federated Learning, Distributed Learning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 26th International Conference on Extending Database Technology; March 25–28, 2025; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learningccs: Computing methodologies Distributed artificial intelligenceccs: Computing methodologies Parallel computing methodologies
@@ 1. Introduction
The big data era has been marked by an unprecedented scale of training datasets (Zhang et al., 2021; Rauf et al., 2024). These datasets are not only growing in size, but are often physically distributed and cannot be easily centralized due to business considerations, privacy concerns, bandwidth limitations (especially in federated settings, such as drones collecting and collaboratively building a global model/view of an area), and data sovereignty laws (Kairouz et al., 2021; Wu et al., 2020; Cormode et al., 2024). Such constraints complicate the use of Deep Learning (DL) techniques in the aforementioned scenarios, and provide challenges in harnessing knowledge from distributed data.
@@
Distributed Deep Learning (DDL) has emerged as an alternative paradigm to the traditional centralized approach (Zinkevich et al., 2010; Chilimbi et al., 2014), offering efficient handling of large-scale data across multiple worker-nodes, enhancing the speed of training DL models and paving the way for more scalable and resilient DL applications (Um et al., 2023; Ma et al., 2023; Lai et al., 2023; Zhou et al., 2022; Davitkova et al., 2024). Most DDL methods are iterative, where, in each iteration, some amount of local training is followed by synchronization of the local models with the global one. The predominant method, based on the bulk synchronous parallel (BSP) approach (Valiant, 1990), is to average the local model updates and then apply the average update to each local model (Zinkevich et al., 2010). Less synchronized variants have also been proposed, to ameliorate the effect of straggler workers (Miao et al., 2021; Fu et al., 2022) but compromise convergence speed and model quality.
@@
A significant challenge inherent in the traditional techniques, especially in federated DL settings, where models are huge and worker interconnections are slow, is the communication bottleneck, restricting system scalability (Tang et al., 2023; Wang et al., 2016). This leads to a low computation-to-communication ratio (Shi et al., 2021; Fu et al., 2022). Addressing the communication overhead to expedite DDL algorithms has been a focal point of research for many years; speeding-up SGD is arguably among the most impactful and transformative problems in machine learning (Wang and Joshi, 2021).
@@
The most direct method to alleviate the communication burden is to reduce the frequency of communication rounds. Local-SGD is the prime example of this approach. It allows workers to perform $\tau$ local update steps on their models before aggregating them, as opposed to averaging the updates in every step (Yu et al., 2019; Haddadpour et al., 2019). Although Local-SGD is effective in reducing communication while maintaining comparable model quality (Wang and Joshi, 2021), determining the optimal value of $\tau$ presents a critical challenge, with only a handful of studies offering theoretical insights into its influence on convergence (Wang and Joshi, 2021; Stich, 2019; Yu et al., 2019).
@@
To further reduce communication costs of Local-SGD, more sophisticated communication strategies introduce varying sequences of local update steps $\{\tau_{0},...,\tau_{R}\}$, instead of a fixed $\tau$. In (Wang and Joshi, [n. d.]), in order to minimize convergence error with respect to wall-time, the authors proposed a decreasing sequence of local update steps. Conversely, the focus in (Haddadpour et al., 2019) was on reducing the number of communication rounds for a fixed number of model updates and an increasing sequence emerged. These contrasting approaches underscore the multifaceted nature of communication strategies in distributed deep learning, highlighting not only the absence of a one-size-fits-all solution but also the growing need for dynamic, context-aware strategies that can continuously adapt to the specific intricacies of the learning task.
@@
Main Idea and Contributions. Our work addresses critical efficiency challenges in DDL, particularly in communication-constrained environments, such as the ones encountered in Federated Learning (FL) applications (Kairouz et al., 2021). We introduce Federated Dynamic Averaging (FDA), a novel, adaptive distributed deep learning strategy that massively improves communication efficiency over previous work.
FDA utilizes a novel 2-action, conditional synchronization protocol, designed to avoid the need to decide or guess the proper values of local update steps, or to synchronize after each training step, but rather only performs the costly synchronization process when needed. Our FDA algorithm dynamically triggers synchronization based on the value of model variance across nodes. In a nutshell, the costly synchronization step is only triggered if the local models have diverged significantly, which implies that the global model may no longer be accurate.
As Figure 1 demonstrates, at the start, nodes enter the local training step with the same global model (Figure 1.(A)). Then, local training commences and each distributed node/worker computes its local state (Figure 1.(B)). This is followed by the transmission (Figure 1.(C)) of these small-size local states, an operation that is bandwidth- and time-efficient because of their small size. During transmission, the local states are aggregated and their average is made available to all nodes—an operation known as AllReduce. This operation does not require (or prohibit) the use of a central node. Based on the aggregated state, the nodes can estimate (Figure 1.(D)) whether the variance of the local models may have exceeded a threshold. If this is not the case, the costly synchronization step (Figure 1.(E)) is avoided and local training continues. What is important is how to properly pick these local states computed at, and then transmitted by, the local nodes. To address this problem, we propose two variants of our FDA algorithm. Our contributions can be summarized as follows:
@@
  • We propose FDA, an algorithm that dynamically decides to synchronize local nodes when model variance across nodes exceeds a threshold. This strategy drastically reduces communication requirements, while preserving cohesive progress towards the shared training objective.
  • We propose two variants of FDA, which differ on the local state that is computed at, and then transmitted by, each distributed node/worker. These two variants are termed SketchFDA and LinearFDA.
@@
  • We evaluate and compare FDA with other DDL algorithms through a comprehensive suite of experiments with diverse datasets, models, and tasks. Our experiments demonstrate that FDA outperforms traditional and contemporary FL algorithms by orders of magnitude in communication, while maintaining equivalent model performance. Furthermore, it effectively balances the competing demands of communication and computation, providing greatly improved trade-offs.
  • We show that FDA remains robust and efficient under various data distributions (Non-IID settings), maintaining comparable performance to the IID case. We note that in Federated Learning scenarios, the norm is that data are not independently and identically distributed across nodes.
Outline. The remainder of this paper is organized as follows: Section 2 reviews related work. Section 3 introduces our DDL technique, Federated Dynamic Averaging (FDA), and its two variants. Section 4 details the experimental setup, and discusses the insights and conclusions drawn from our empirical investigation. Lastly, Section 5 contains concluding remarks.
Figure 1. FDA. The local training step is followed by the computation of a local state by all nodes. Then, the (small in size) local states are aggregated. Based on the aggregated result, all nodes estimate if synchronization is required. In most cases, the expensive synchronization step of the models is avoided and local training continues
@@ 2. Related Work
where $K$ is the number of workers and $F_{k}(\mathbf{w})\triangleq\mathbb{E}_{\zeta_{k}\sim\mathcal{D}_{k}}\left[\ell(\mathbf{w};\zeta_{k})\right]$ is the local objective function for worker $k$. Function $\ell(\mathbf{w};\zeta_{k})$ represents the loss for data sample $\zeta_{k}$ given model $\mathbf{w}$.
@@
Communication efficient Local-SGD. The work in (Li et al., 2021) decomposes each round into two phases. In the first phase, each worker runs Local-SGD with $\tau=I_{1}$, while the second phase run $I_{2}$ steps with $\tau=1$(Li et al., 2021) proposes to exponentially decay $I_{1}$ every $M$ rounds. In the heterogeneous setting, the work in (Qin et al., 2023), by analysing the convergence rate, proposes an increasing sequence of local update steps for strongly-convex local objectives, and suggests fixed local steps for convex, non-convex local objectives. The study in (Yu and Jin, 2019) dynamically increases batch sizes to reduce communication rounds, maintaining the same convergence rate as SSP-SGD. However, the large-batch approach leads to poor generalization (Hoffer et al., 2017), a challenge addressed by the post-local SGD method (Lin et al., 2020), which divides training into two phases: BSP-SGD followed by Local-SGD with a fixed number of steps. In the Lazily Aggregated Algorithm (LAG) (Chen et al., 2018), a different approach was taken, using only new gradients from some selected workers and reusing the outdated gradients from the rest, which essentially skips communication rounds.
@@
Federated Averaging (FedAvg) (McMahan et al., 2017) is another representative communication efficient Local-SGD algorithm, which is a pivotal method in Federated Learning (FL) (Kairouz et al., 2021). In the FL setting with edge computing systems, the work in (Wang et al., 2019) tries to find the optimal synchronization period $\tau$ subject to local computation and aggregation constraints. Recently (Mills et al., 2023), in the FL setting with the assumption of strongly-convex objectives, by analysing the balance between fast convergence and higher-round completion rate, a decaying local update step scheme emerged.
@@
Accelerating convergence. An indirect, yet highly effective, way to mitigate the communication burden in DDL, is to speed up convergence. Consequently, recent works have deployed accelerated versions of SGD to the distributed setting. Specifically, FedAdam (Reddi et al., 2021) extends Adam (Kingma and Ba, 2015) and FedAvgM (Hsu et al., 2019) extends SGD with momentum (SGD-M) (Sutskever et al., 2013a). Recently, Mime (Karimireddy et al., 2021) provides a framework to adapt arbitrary centralized optimization algorithms to the FL setting. However, these methods still suffer from the model divergence problem, particularly in heterogeneous settings. When solving (1), the disparity between each worker’s optimal solution $\mathbf{w}_{k}^{*}$ for their objective $F_{k}$, and the global optimum $\mathbf{w}^{*}$ for $F$, can potentially cause worker models to diverge (drift) towards their disparate minima (Karimireddy et al., 2020; Reddi et al., 2021; Wenig and Papenbrock, 2022). The result is slow and unstable convergence with significant communication overhead. To address this problem, the SCAFFOLD algorithm (Karimireddy et al., 2020) used control-variates (in the same spirit to SVRG), with significant speed-up. FedProx (Sahu et al., 2018) re-parameterized FedAvg (McMahan et al., 2017) by adding $L^{2}$ regularization in the workers’ objectives to be near the global model. Lastly, FedDyn (Acar et al., 2021) improved upon these ideas with a dynamic regularizer making sure that if local models converge to a consensus, this consensus point aligns with the stationary point of the global objective function.
@@
Compression. To reduce communication overhead in DDL, significant efforts have been directed towards minimizing message sizes. Key strategies include sparsification, where only crucial components of information are transmitted, as explored in (Aji and Heafield, 2017), and quantization techniques, which involve transmitting only quantized gradients, as detailed in (Shlezinger et al., 2021). Notably, these compression techniques can be combined with Local-SGD methods to enhance communication-efficiency further. A prime example is Qsparse-local-SGD (Basu et al., 2019), which integrates aggressive sparsification and quantization with Local-SGD, achieving substantial communication savings to reach specific target accuracies.
Additionally, sketching emerges as another fundamental tool in large-scale machine learning. It effectively compresses high-dimensional problems into lower dimensions to save runtime and memory, typically utilizing hash-based probabilistic data structures. For instance, (Spring et al., 2019) use Count Sketches to compress auxiliary variables in optimization algorithms, significantly freeing up memory. Similarly, FetchSGD (Rothchild et al., 2020) employs Count Sketches to compress model updates and leverages their linearity for efficient merging. In contrast to these applications, our approach utilizes sketches not for compression but to estimate local state information, and based on this to decide whether a synchronization is required—an orthogonal application to traditional use cases. A comprehensive survey of compression techniques in DDL can be found in (Wang et al., 2023).
@@ 3. Federated Dynamic Averaging
We now present our algorithms, based on our notion of Federated Dynamic Averaging (FDA). Our algorithms deviate from prior work in these two key ways:
  1. (1)
    The decision on when to synchronize.
  2. (2)
    The actual synchronization process.
To the best of our knowledge, this is the first Distributed Deep Learning algorithm that dynamically decides when to synchronize based on the current collective state of the training progress—whether it is advancing well or poorly.
@@
Notation. At each time step $t$, each worker $k$ independently maintains its own vector of model parameters [^1] The terms “model” and ”model parameters” are used interchangeably, as is common in the literature., denoted as $\mathbf{w}_{t}^{(k)}\in\mathbb{R}^{d}$. Let $\mathbf{w}_{t}$ represent the $K\times d$ tensor of all local model vectors, and $\overline{\mathbf{w}}_{t}$ be the average model vector (this notation applies to all vector quantities):
$\mathbf{w}_{t}=\left[\mathbf{w}_{t}^{(1)},\ldots,\mathbf{w}_{t}^{(K)}\right]\quad,\quad\overline{\mathbf{w}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{t}^{(k)}$
Furthermore, let $\textsc{Optimize}(\mathbf{w},\mathcal{B})$ be the updated model vector (Goodfellow et al., 2016) computed by some optimization algorithm (e.g., SGD, Adam) using the model $\mathbf{w}$, and the batch $\mathcal{B}$ of training data. It incorporates the learning rate, loss function and relevant gradients. During step $t$, each worker $k$ first applies the update:
$\mathbf{w}^{(k)}_{t}=\textsc{Optimize}(\mathbf{w}_{t-1}^{(k)}\,,\,\mathcal{B}_{t}^{(k)})$
Moreover, operation $\textsc{AllReduce}(\mathbf{w}_{t}^{(k)})$ computes and returns the average model vector (Li et al., 2020):
$\overline{\mathbf{w}}_{t}=\textsc{AllReduce}(\mathbf{w}_{t}^{(k)})$
In prior work, workers may (or not) synchronize by executing $\textsc{AllReduce}(\mathbf{w}_{t}^{(k)})$, thereby setting $\mathbf{w}_{t}^{(k)}=\overline{\mathbf{w}}_{t}$. If synchronization is not performed at step $t$, then each worker continues with the locally updated model.
Table 1. Notation
Symbol Meaning
[…]
$t$ Time step index
$K$ Number of workers
$d$ Model dimension
$\mathbf{w}_{t}^{(k)}\in\mathbb{R}^{d}$ Model of worker $k$
$\mathbf{w}_{t}=[\mathbf{w}_{t}^{(1)},\ldots,\mathbf{w}_{t}^{(K)}]$ Tensor of local models
$\overline{\mathbf{w}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{t}^{(k)}$ Average model vector
$\overline{\mathbf{w}}_{t_{0}}$ Model after most recent sync.
$\mathbf{u}_{t}^{(k)}=\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t_{0}}$ Local model drift
@@
$\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ Model variance
$\Theta$ Model variance threshold
$\mathbf{\sigma}_{t}^{(k)}$ Lower-dim. proxy of $\mathbf{u}_{t}^{(k)}$
$\mathbf{S}_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\mathbf{\sigma}_{t}^{(k)}\right)$ State of worker $k$
$\overline{\mathbf{S}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{S}_{t}^{(k)}$ Average state
Model Variance and FDA. The model variance quantifies the dispersion or spread of worker models around the average model:
@@
(2) $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t}\right\|_{2}^{2}$
@@
This measure provides insight into how closely aligned the workers’ models are at any given time. High variance indicates that the models are widely spread out, essentially drifting apart, leading to a lack of cohesion in the aggregated model. Conversely, a moderate or low variance suggests that the workers’ models are closely aligned, working collectively towards the shared objective.
@@
The FDA algorithm (Algorithm 1) is based on the premise that, as long as the variance is below a threshold $\Theta$, synchronization is not needed. Thus, we introduce the Round Invariant (RI):
@@
To preserve the RI, our FDA algorithm maintains (Lines 4-6 of Algorithm 1) at each node $k$ a local (low-dimensional) state-vector $\mathbf{S}_{t}^{(k)}$, which is computed based on $\mathbf{w}_{t}^{(k)}$. These state vectors are vital for the subsequent estimation of the model variance, and underpin the two variants of the FDA algorithm (provided in Sections 3.1 and 3.2, respectively). Our estimation techniques begin by performing AllReduce on the states $\mathbf{S}_{t}^{(k)}$, consolidating them into the average state $\overline{\mathbf{S}}_{t}$ (Line 7). Importantly, this communication step requires significantly less bandwidth and resources than transmitting the full models $\mathbf{w}^{(k)}_{t}$.
For each FDA variant, we also define a (different) function $H(\overline{\mathbf{S}}_{t})$ that overestimates the variance, i.e., it ensures that as long as $H(\overline{\mathbf{S}}_{t})\leq\Theta$ then the variance is bounded by $\Theta$. This guarantee is probabilistic for the Sketch-based variant of FDA, and deterministic for its Linear counterpart. Consequently, if $H(\overline{\mathbf{S}}_{t})>\Theta$ then synchronization is performed (Lines 8-9) — the RI invariant cannot be guaranteed. After synchronization, the model variance is zero.
Efficiently Monitoring the RI. Estimating model variance efficiently is at the heart of FDA. To this end, we first introduce the local model drift
$\mathbf{u}_{t}^{(k)}=\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t_{0}}$
where $\overline{\mathbf{w}}_{t_{0}}$ is the model vector after the most recent synchronization. Subsequently, the model variance can be written as: